OpenStudy (anonymous):

Which of the following is a factor of x2 + 19x + 84? x + 6 x + 7 x + 2 None of the above

OpenStudy (callisto):

x^2 + 19x + 84 = x^2 + 7x + 12x + 84 = x(x+7) + 12(x+7) = ...?!

Parth (parthkohli):

You may use the factor theorem too here. Check which of the following results in 0. 1) x = -6 2) x = -7 3) x = -2 4) None

OpenStudy (anonymous):

x=7 ?

Parth (parthkohli):

Don't get confused by my answer. Check out @Callisto 's

OpenStudy (anonymous):

i dont understand